OpenStudy (anonymous):

Rewrite the complex number -3 +4i into trigonometric form. Give the argument, theta, to the nearest degree.

OpenStudy (earthcitizen):

\[z = \cos(\theta) +isin(\theta)\]use De Moivre's theorem, where z = x+iy first find the argument and modulus of the complex number and the angle then plug it into

OpenStudy (earthcitizen):

\[r = 5, \theta = 126.87^o\]\[z = r(\cos(\theta) + isin(\theta))\]